OpenStudy (anonymous):

change each percent to a fraction. 85%, 2.8%

OpenStudy (anonymous):

2 and 4 fiths for 2.8

OpenStudy (anonymous):

\[85/100=17/20\] 2.8/100=28/1000=7/250